### v1

@@ Results
Finally, we tested the hierarchical annealing process in random networks [33], where the probability of two nodes being connected is uniform. The topology of Erdős-Rényi networks has been well characterized and is considered a valid null system for multiple graphs and topological metrics in biological systems. Again, our solution produced comparably optimal solutions for all algorithms (Fig. 2c).
Figure 2: Annealing solutions for different random networks of increasing size. Modularity of a random network as found by the Louvain, Leiden, and Hierarchical annealing algorithms. The resolution, $\gamma$, was set to 1. To measure the overlap between the different solutions, we show the Dice score between all the communities found. The colormaps indicate if the annealing solution has a higher (in blue), equal (in black), or lower (in red) modularity. The opacity is directly proportional to the magnitude of the Dice between each pair of communities. The random networks have been generated via the mechanisms of (a) power-law clustering, (b) preferential attachment, and (c) random edge creation.
@@
%$\%RI\sim 1+\alpha+N.$
This model carried explanatory power ($R^{2}=0.3680$, $p=0.0080$, $F$-statistic against a constant model, AIC=118.2572) and both coefficients significantly contributed to the prediction ($\alpha$: $p=0.0.0116$, $N$: $p=0.0462$). Additionally, we included an interaction term of the form $\alpha N$. This alternative model showed a slightly higher coefficient of determination ($R^{2}=0.3869$, $p=0.0144$, $F$-statistic against a constant model, AIC=119.5288) but the individual coefficients were non-significant ($\alpha$: $p=0.0863$, $N$: $p=0.7353$, $\alpha N$: $p=0.4416$), suggesting that independent rather than interaction effects better explain the performance of the algorithm.
@@ Discussion
Biological and social networks are dynamic rather than static systems, even if we study their properties at a fixed point in time. Their evolution is thus constrained to the underlying mechanism that governs the interactions between their constituents. Barabasi and Albert [34] showed that preferential attachment, where highly connected nodes attract more links, leads to scale-free networks with a slower decay of highly connected nodes than in random networks. Preferential attachment can lead to structural hierarchies, but the latter is not guaranteed [35, 39]. Furthermore, life processes themselves are hierarchical [40, 41], inspiring specific methods to discover such organization [42]. In neuroscience, brain network structure is influenced by plasticity mechanisms occurring across temporal and spatial scales. Hebbian plasticity, in particular, has been shown to modulate preferential attachment, resulting in scale-free and clustered network structures [43]. It is therefore not unreasonable to assume some underlying hierarchical structure to design the algorithms. In the context of our work, we show how the Hierarchical annealing process performs exquisitely in these kinds of systems even improving over classical alternatives (Fig. 3). Additionally, network size did not significantly affect performance, suggesting the process is robust and scalable. The slightly suboptimal results we observed likely stem from the hierarchical process itself, rather than issues with the annealing steps [44]. Crucially, our results seem to indicate the latter, as optimal solutions are returned for multiple types of networks.
@@
The specific case where the resolution parameter $\gamma$ is 1, for which we recover known results in the literature [3], simplifies several expressions throughout this work. However, discovering communities via maximizing the modularity function possesses an inherent limitation where communities smaller than a given size are likely to be omitted [27]. Kumpula, et al., proved that this limit persisted for all $\gamma$ independently of the null model used in the definition of $Q$ [45]. Although the resolution parameter was introduced to weigh the topological scale at which communities ought to be detected [4], it was elegantly shown that it could be employed to overcome the resolution limit [46], albeit disregarding the null model (e.g., $\frac{g_{i}g_{j}}{2m}$). Consequently, approaches that use different definitions for modularity function and the parameter $\gamma$ also exist [47, 48, 44, 49], but they still fall within spin-glass systems, therefore being possible to tackle using the formalism we used.
@@
Nonetheless, Newman and Girvan’s quality index [1] along with the explicit inclusion of a resolution parameter [4], remains, up to date, the most common choice for discovering community structure but, unfortunately, the exact value remains, up to date, an arbitrary choice. In short, large values of the resolution will yield smaller communities, while smaller values of $\gamma$ will result in fewer communities of increased size. Hence, we root the parameter $\gamma$ inside our framework for two main reasons: 1) users need not tune the objective function to maximize, and, 2) it easily adapts to multiresolution methods. We maintain $\gamma$ as a free parameter and we derive all our expressions without specifying its value. Our experiments using multiple values of $\gamma$ within a meaningful range show that the Hierarchical annealing process is robust w.r.t. the chosen resolution; even in networks without inherent hierarchies (Fig. S1). This feature allows us to speculate that QA can be effectively used to detect community structure using multiresolution methods, where the resolution is gradually changed or sampled nonuniformly along the optimization process [50]. Further work should address whether this holds or not, the range for which the annealer cannot coordinate the large number of communities found, and the computational time required to obtain the large number of relatively small-size communities expected for $\gamma>3$.
@@
A limitation of this work is the increased computational time compared to classical heuristics. However, simulated annealing returns very good results at the cost of increased computational times [5]. Our preliminary experiments suggest a roughly linear relationship between the number of communities detected and the elapsed time. Some observed irregularities in timing may be due to repeated instances and calls to the QPUs. Although we mitigated this by caching precomputed graph embeddings, we were unable to fully address the causes of these longer times. While total times never exceeded one minute per network per run, this remains a limitation for larger, highly segregated networks. Notably, the average time consumed by the QPU per binary division, as reported by the D-Wave portal, was approximately 2 milliseconds, comparable to times for the Louvain and Leiden methods. Thus, with further research, quantum annealing and its combination with algorithmic processes could potentially become competitive in this regard.
@@ Conclusions
In this work, we use the D-Wave Advantage annealer to maximize the modularity function to discover community structure in complex networks. We demonstrate that hierarchically dividing the network and, at each step, optimizing a binary problem effectively eliminates the need to use one-hot encoding and constraints to maximize Newman’s quality function. As opposed to earlier attempts, the number of communities can be left uncontrolled and found by the algorithm. Moreover, extending Newman’s ideas [3, 28], we analytically show how the function to maximize remains formally independent from the resolution parameter and whether the network is undirected, directed, and/or weighted. We thoroughly conduct experiments to assess whether the underlying topology of the network represents a liability to this nested approach finding that the pure quantum process is stable across topologies and even performs better in scale-free networks, which are ubiquitous in biological systems. Furthermore, the same procedure could, in principle, be applied for multiresolution clustering purposes in any given type of network given that the Hierarchical annealing optimization remains largely equivalent to classical counterparts for multiple values of the resolution parameter.
@@ Acknowledgements
The publication was created within the project of the Minister of Science and Higher Education ”Support for the activity of Centers of Excellence established in Poland under Horizon 2020” on the basis of the contract number MEiN/2023/DIR/3796. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 857533. This publication is supported by Sano project carried out within the International Research Agendas programme of the Foundation for Polish Science, co-financed by the European Union under the European Regional Development Fund. The research presented in this paper received support from the funds assigned by Polish Ministry of Science and Technology to AGH University. We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2024/017319.
@@
J.F.-R. and K.R. conceptualized, designed, and supervised research; J.F.-R., K.J., B.W., and K.R. designed the software package; K.J. and B.W. implemented the software package; J.F.-R. designed and performed the experiments with the help of K.J. and B.W.; J.F.-R., K.C., and K.R. analyzed the results; J.F.-R. created the figures and wrote the first draft of the manuscript; all authors revised, edited, and approved the final version of the manuscript.
@@
Joan Falcó-Roget1*, Kacper Jurek2,3, Barbara Wojtarowicz2, Katarzyna Rycerz2,3

### v2

@@ Results
We also tested the hierarchical annealing process in random networks [33], where the probability of two nodes being connected is uniform. The topology of Erdős-Rényi networks has been well characterized and is considered a valid null system for multiple graphs and topological metrics in biological systems. Again, our solution produced comparably optimal solutions for all algorithms (Fig. 2c).
Finally, we studied the performance of the hierarchical quantum annealing solution in directed scale-free networks generated using the preferential attachment mechanism [36]. Again, our solution resulted in high-quality community structures for multiple network sizes (Fig. 2d), and even in important increments in certain cases w.r.t. the classical benchmarks (Fig. 2e).
Figure 2: Annealing solutions for different random networks of increasing size. Modularity of a random network as found by the Louvain, Leiden, and Hierarchical annealing algorithms. The resolution, $\gamma$, was set to 1. To measure the overlap between the different solutions, we show the Dice score between all the communities found. The colormaps indicate if the annealing solution has a higher (in blue), equal (in black), or lower (in red) modularity. The opacity is directly proportional to the magnitude of the Dice score between each pair of communities. The random networks have been generated via the mechanisms of (a) power-law clustering, (b) preferential attachment, and (c) random edge creation. d Exact same layout for a directed scale-free network. e Community structure of a directed scale-free network of 60 nodes discovered through the H. annealing and Louvain algorithms. This corresponds to the network where the annealing solution was significantly superior in d.
@@
%$\%RI\sim 1+\alpha+N,$
where $\alpha$ denotes the power-law exponent, and $N$ is the number of nodes. This model carried explanatory power ($R^{2}=0.3680$, $p=0.0080$, $F$-statistic against a constant model, AIC=118.2572) and both coefficients significantly contributed to the prediction ($\alpha$: $p=0.0.0116$, $N$: $p=0.0462$). Additionally, we included an interaction term of the form $\alpha N$. This alternative model showed a slightly higher coefficient of determination ($R^{2}=0.3869$, $p=0.0144$, $F$-statistic against a constant model, AIC=119.5288) but the individual coefficients were non-significant ($\alpha$: $p=0.0863$, $N$: $p=0.7353$, $\alpha N$: $p=0.4416$), suggesting that independent rather than interaction effects better explain the performance of the algorithm.
@@
Application of QA to structural brain connectivity
One of the main advantages of our proposal is that we could inspect the output process at every step of the process. We plotted the modularity and the community structure in the form of a dendrogram. The final division was dependent on each earlier subdivision, thus being interpretable as a hidden community structure. Such hidden structures would be informative of certain network vulnerabilities or strengths. We applied this visualization technique to all the networks studied up until now, while comparing the modularity of the final community structure with the same benchmark methods achieving competitive and identical results (Fig. S.3; see also Supplementary Figures).
Figure 5: Dendrogram of a clustered power-law network [35] of N=100 nodes. Hierarchical structure found by the H. annealing algorithm. The hierarchy corresponds to the one unraveled after maximizing the modularity of the network as described in the main text; that is, num_runs = 15 and resolution = 1. Both the Louvain and Leiden algorithms returned a maximum modularity of $Q=0.7779$. The left axis shows the modularity at each step of the process, while the right axis displays the corresponding increments. The bottom-most row represents the final output of the H. annealing algorithm.
As a last proof of concept, we analyzed real brain connectivity networks as derived from diffusion-weighted magnetic resonance imaging methods. These networks contained a total of 166 nodes and were densely connected while exhibiting somewhat scale-free structures [40]. We also inspected the anatomical positioning and organization of the found community structure as well as the hierarchical levels of the process (Fig. 6). Despite the large number of variables (i.e., qubits), the H. annealing solution was strikingly close to the solutions found by the Louvain and Leiden algorithms. Furthermore, the communities found by the quantum algorithm only differed slightly in the frontoparietal border, areas that are highly interconnected with U-fibers. The rest of the identified communities were largely in agreement between algorithms (see Supplementary Figures), hence setting a very promising set of baseline results for more advanced and sophisticated QA methods and algorithms.
Figure 6: Dendrogram of a real brain structural connectivity network [40]. a Hierarchical structure found by the H. annealing algorithm. The hierarchy corresponds to the one unraveled after maximizing the modularity of the network as described in the main text; that is, num_runs = 10 and resolution = 1. The Louvain and Leiden algorithms returned a maximum modularity of $Q=0.611$ and $Q=0.610$ respectively. The left axis shows the modularity at each step of the process, while the right axis displays the corresponding increments (i.e., Eq. 9). The bottom-most row represents the final output of the H. annealing algorithm. b Overlay of the structural communities on the Montreal Neurological Institute (MNI) 152 template. Each color corresponds to a different community, whose elements can be found in a.
@@ Discussion
Unveiling hierarchies in complex networks with quantum computing
@@
Biological and social networks are dynamic rather than static systems, even if we study their properties at a fixed point in time. Their evolution is thus constrained to the underlying mechanism that governs the interactions between their constituents. Barabasi and Albert [34] showed that preferential attachment, where highly connected nodes attract more links, leads to scale-free networks with a slower decay of highly connected nodes than in random networks. Preferential attachment can lead to structural hierarchies, but the latter is not guaranteed [35, 42]. Furthermore, life processes themselves are hierarchical [43, 44], inspiring specific methods to discover such organization [45]. In neuroscience, brain network structure is influenced by plasticity mechanisms occurring across temporal and spatial scales. Hebbian plasticity, in particular, has been shown to modulate preferential attachment, resulting in scale-free and clustered network structures [46]. It is therefore not unreasonable to assume some underlying hierarchical structure to design the algorithms. In the context of our work, we show how the Hierarchical annealing process performs exquisitely in these kinds of systems even improving over classical alternatives (Fig. 3). Additionally, network size did not significantly affect performance, suggesting the process is robust and scalable. The slightly suboptimal results we observed likely stem from the hierarchical process itself [47], rather than issues with the annealing steps given that we mitigated its variability by using multiple and independent instances of the quantum hierarchical process.
One of the main advantages of the hierarchical approach we studied, is the direct interpretability of each one of the recursive divisions. We can construct dendrograms of the sequence of divisions that yield the highest modularity indices for any given network. However, the correctness of such dendrograms depends entirely on the optimality of the represented solution. For that, it remains essential to compare the modularity of the community structure with solutions from other algorithms. As expected, the obtained using H. annealing for all the networks studied in this work were extremely competitive, thus offering insights into the hidden structure of these networks (Figs. S.3 and 6; see also Supplementary Figures). There exist some attempts to reveal hierarchies present in complex networks, most notably relying on bottom-up hierarchical clustering methods [45, 47] and the combination of resolution parameters [48], therefore offering a complementary rather alternative description of fractal-like structure. A few other methods have been developed using the Louvain algorithm to unravel hierarchical relationships within complex systems. For example, each iteration of the Louvain algorithm can be interpreted as a step in a hierarchical process [10, 49]. Alternatively, separate instances of the Louvain algorithm may be used to discover community structures within sub-modules [50]. Instead, our approach considers the whole system at each step inside the hierarchy [3, 28] (see also Eq. 9). Because of that, and contrary to the Louvain solution, this guarantees that every step in the hierarchy is divided taking into account the whole system rather than local sub-modules, a desirable feature for strongly nested systems where our method is expected to outperform several alternatives (Fig. 3; see also Supplementary Figures).
Towards quantum-based multi-resolution community detection methods
@@
The specific case where the resolution parameter $\gamma$ is 1, for which we recover known results in the literature [3], simplifies several expressions throughout this work. However, discovering communities via maximizing the modularity function possesses an inherent limitation where communities smaller than a given size are likely to be omitted [27]. Kumpula, et al., proved that this limit persisted for all $\gamma$ independently of the null model used in the definition of $Q$ [51]. Although the resolution parameter was introduced to weigh the topological scale at which communities ought to be detected [4], it was elegantly shown that it could be employed to overcome the resolution limit [52], albeit disregarding the null model (e.g., $\frac{g_{i}g_{j}}{2m}$). Consequently, approaches that use different definitions for modularity function and the parameter $\gamma$ also exist [53, 54, 47, 55]. However, they still fall within spin-glass systems, therefore being possible to tackle using the same formalism we used.
@@
Nonetheless, Newman and Girvan’s quality index [1] along with the explicit inclusion of a resolution parameter [4], remains, up to date, the most common choice for discovering community structure but, unfortunately, the exact value remains, up to date, an arbitrary choice. In short, large values of the resolution will yield smaller communities, while smaller values of $\gamma$ will result in fewer communities of increased size. Hence, we root the parameter $\gamma$ inside our framework for two main reasons: 1) users need not tune the objective function to maximize, and, 2) it easily adapts to multiresolution methods. We maintain $\gamma$ as a free parameter and we derive all our expressions without specifying its value. Our experiments using multiple values of $\gamma$ within a meaningful range show that the Hierarchical annealing process is robust w.r.t. the chosen resolution; even in networks without inherent hierarchies (Fig. S1). This feature allows us to speculate that QA can effectively detect community structure using multiresolution methods, where $\gamma$ is gradually changed or sampled nonuniformly along the optimization process [56]. Further work should address whether this holds or not, the range for which the annealer cannot coordinate the large number of communities found, and the computational time required to obtain the large number of relatively small-size communities expected for $\gamma>3$.
@@
Limitations and future work
A limitation of this work is the increased computational time compared to classical heuristics. However, simulated annealing returns very good results at the cost of increased computational times [5]. Our preliminary experiments suggest a roughly linear relationship between the number of communities detected and the elapsed time. Some observed irregularities in timing may be due to repeated instances and calls to the QPUs. Although we mitigated this by caching precomputed graph embeddings, we were unable to fully address the causes of these longer times. While total times never exceeded one minute per network per run, this remains a limitation for larger, highly segregated networks. Notably, the average time consumed by the QPU per binary division, as reported by the D-Wave portal, was approximately 2 milliseconds, comparable to times for the Louvain and Leiden methods. Thus, with further research, quantum annealing and its combination with algorithmic processes could potentially become competitive in this regard while increasing its accuracy.
@@ Conclusions
In this work, we use the D-Wave Advantage annealer to maximize the modularity function to discover community structure in complex networks. We demonstrate that hierarchically dividing the network and, at each step, optimizing a binary problem effectively eliminates the need to use one-hot encoding and constraints to maximize Newman’s quality function. As opposed to earlier attempts, the number of communities can be left uncontrolled and found by the algorithm. Moreover, extending Newman’s ideas [3, 28], we analytically show how the function to maximize remains formally independent from the resolution parameter and whether the network is undirected, directed, and/or weighted. We thoroughly conduct experiments to assess whether the underlying topology of the network represents a liability to this nested approach finding that the pure quantum process is stable across topologies and even performs better in scale-free networks, which are ubiquitous in biological systems. Furthermore, the same procedure could, in principle, be applied for multiresolution clustering purposes in any given type of network given that the Hierarchical annealing optimization remains largely equivalent to classical counterparts for multiple values of the resolution parameter. Lastly, the hierarchical annealing process we present allows for interpretable results. When adequately compared to classical alternatives, it permits the exploration of hidden hierarchies within complex systems, potentially indicative of several network dysfunctions.
@@ Acknowledgements
The publication was created within the project of the Minister of Science and Higher Education ”Support for the activity of Centers of Excellence established in Poland under Horizon 2020” on the basis of the contract number MEiN/2023/DIR/3796. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 857533. This publication is supported by Sano project carried out within the International Research Agendas programme of the Foundation for Polish Science, co-financed by the European Union under the European Regional Development Fund (J.F.-R., B.W., K.C., and K.R.). The research presented in this paper received support from the funds assigned by Polish Ministry of Science and Technology to AGH University (K.J., K.C., and K.R.). We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2024/017739 (J.F.-R., K.J., B.W., and K.R.).
@@
J.F.-R. and K.R. conceptualized, designed, and supervised research; J.F.-R., K.J., B.W., and K.R. designed the software package; K.J. and B.W. implemented the software package and the code to visualize the dendrograms; J.F.-R. designed and performed the experiments with the help of K.J. and B.W.; J.F.-R., K.C., and K.R. analyzed the results; J.F.-R. created the figures and wrote the first draft of the manuscript; all authors revised, edited, and approved the final version of the manuscript.
@@
Joan Falcó-Roget1*, Kacper Jurek2,3, Barbara Wojtarowicz1,2, Karol Capała1,2, Katarzyna Rycerz1,2,3
@@ 3 Quantum Computing Lab, Academic Computer Center Cyfronet AGH, Nawojki 11, Krakow, 30-950, Poland
Figure S.2: Dendrogram of the Karate network [39]. Hierarchical structure found by the H. annealing algorithm. The hierarchy corresponds to the one unraveled after maximizing the modularity of the network as described in the main text; that is, num_runs = 50 and resolution = 1. The left axis shows the modularity at each step of the process, while the right axis displays the corresponding increments. The bottom-most row represents the final output of the H. annealing algorithm.
Figure S.3: Dendrogram of a clustered power-law network [35] of N=80 nodes. Hierarchical structure found by the H. annealing algorithm. The hierarchy corresponds to the one unraveled after maximizing the modularity of the network as described in the main text; that is, num_runs = 10 and resolution = 1. Both the Louvain and Leiden algorithms returned a maximum modularity of $Q=0.7563$. The left axis shows the modularity at each step of the process, while the right axis displays the corresponding increments. The bottom-most row represents the final output of the H. annealing algorithm.
Figure S.4: Dendrogram of a Barabasi-Albert network [34] of N=70 nodes. Hierarchical structure found by the H. annealing algorithm. The hierarchy corresponds to the one unraveled after maximizing the modularity of the network as described in the main text; that is, num_runs = 20 and resolution = 1. The Louvain and Leiden algorithms returned maximum modularities of $Q=0.0760$ and $Q=0.0698$ respectively. The left axis shows the modularity at each step of the process, while the right axis displays the corresponding increments. The bottom-most row represents the final output of the H. annealing algorithm.
Figure S.5: Dendrogram of an Erdos-Renyi network [33] of N=70 nodes. Hierarchical structure found by the H. annealing algorithm. The hierarchy corresponds to the one unraveled after maximizing the modularity of the network as described in the main text; that is, num_runs = 50 and resolution = 1. The Louvain and Leiden algorithms returned maximum modularities of $Q=0.1189$ and $Q=0.1021$ respectively. The left axis shows the modularity at each step of the process, while the right axis displays the corresponding increments. The bottom-most row represents the final output of the H. annealing algorithm.
Figure S.6: Dendrogram of a directed sclae-free network [36] of N=50 nodes Hierarchical structure found by the H. annealing algorithm. The hierarchy corresponds to the one unraveled after maximizing the modularity of the network as described in the main text; that is, num_runs = 50 and resolution = 1. Both the Louvain and Leiden algorithms returned a maximum modularity of $Q=0.31975$.
Figure S.7: Structural communities found at the early steps of the H. annealing process. We plot the community structure discovered during the 2nd division of the hierarchical method described in the main text. It is visible how the hierarchy carries concise anatomical information, thus being a valid way to inspect hidden structures within complex networks through quantum annealing optimization.
Figure S.8: Structural communities found by the Leiden and Louvain algorithms. The resemblance between these communities and the ones found by the annealing process is strikingly high (see main text). The modularity found by the Leiden algorithm was $Q=0.608$, and the modularity returned by the Louvain algorithm was $Q=0.611$.
Software specifics: Qommunity
Architecture
Figure S.9: Architecture of Qommunity library. The package makes use of the quantum computing library QHyper [65] to communicate with the quantum annealer using multiple instances of a quadratic binary optimization problem. The Qommunity library also allows the user to choose from several other algorithms and solvers described in the main text. Lastly, the modularity maximization happens at the IterativeSearcher class, where any solver and algorithm can be explicitly and automatically run an arbitrary number of times to produce a sufficiently big number of solutions to choose from.
The Qommunity library was designed to integrate several methods for detecting communities in graphs in one place. The implemented methods rely on classical, quantum, and hybrid solutions. The idea was to simplify and standardize the interface for conducting experiments. Several components compartmentalize the architecture, offering user-friendly features: Samplers, Searchers, and IterativeSearchers (Fig. S.9).
Samplers: Through Samplers we specify the method by which we want to detect communities, providing the necessary parameters. Currently, several classical solutions are implemented, such as Gurobi, Louvain [10], Bayan[15], Leiden [12], hybrid – using DQM [25], and purely quantum – using Advantage [19, 65]. We divided samplers based on how the divisions are done;
  • regular: the dividing into communities is done according to the algorithms specified inside the Sampler.
  • hierarchical: perform binary divisions using classical methods or quantum annealing, depending on the Sampler specified. If no division is optimal, the Sampler returns a single community. Alternatively, each binary division can be further continued in the course of recursion logic provided by the HierarchicalSearcher.
Although Samplers provide community detection methods and can perform the community divisions themselves, the Seacher and IterativeSearcher are the dedicated tools for the user to perform the community detection process comprehensively.
Searchers: Searchers are responsible for executing community detection and for the appropriate postprocessing of its results. There are two types of Searchers - hierarchical, which accept Samplers that perform hierarchical divisions of communities, and regular, which perform a non-hierarchical, regular community detection method. Each option is defined by the algorithm of choice as follows:
  • regular: they serve as wrappers for the RegularSampler, facilitating the regular community detection.
  • hierarchical: they extend the scope of action far beyond the RegularSearchers, by implementing the logic of recursion in the hierarchical search. They build the construct of recursive calls. Its execution results in a tree of binary splits. Similar to RegularSearchers, they also wrap the appropriate Sampler, but with a different purpose - to call the community detection method on HierarchicalSampler on each recursion level and, depending on whether a binary split has occurred as its result, decide whether to proceed with recursion. Hence, they build and oversee the process of hierarchical community detection search, described in Algorithm 1 in the Main text.
    The hierarchical searchers track the trace of the hierarchical divisions and, if specified by the user, can return it as a division tree, alongside division modularities, corresponding to each hierarchy division level.
Iterative Searchers: Iterative Searchers serve as wrappers for community Searchers, responsible for calling their corresponding (i.e. regular or hierarchical) community search method a given number of times. These multiple calls are useful due to the stochastic and probabilistic natures of the algorithms Qommunity provides. As a result, for each iteration, we get a graph divided into communities, its modularity measure, and the execution time. The result may also contain the division trees and the division modularities, in the case of hierarchical community search. Iterative Searchers are used to simplify the execution of experiments aimed at maximizing the modularity of a given network (see Algorithm 2 in the Main text).
High-level Python code examples to use the quantum resources for modularity maximization and interpretation
1import networkx as nx # Hagberg, et al. 2008
2import numpy as np
3
4## ORIGINAL FROM THIS WORK
5from Qommunity.samplers.hierarchical.advantage_sampler import AdvantageSampler # H. Annealing
6from Qommunity.samplers.hierarchical.gurobi_sampler import GurobiSampler # H. Gurobi
7
8## HYBRID QUANTUM SOLVER (Wierzbinski, et al. 2023)
9from Qommunity.samplers.regular.dqm_sampler import DQMSampler
10
11## ALETERNATIVE BENCHAMRKS
12from Qommunity.samplers.regular.bayan_sampler import BayanSampler
13from Qommunity.samplers.regular.leiden_sampler import LeidenSampler
14from Qommunity.samplers.regular.louvain_sampler import LouvainSampler
15
16## MODULARITY MAXIMIZATION
17from iterative_searcher.iterative_searcher import IterativeSearcher
18
19## NETWORK TO ANALYZE
20Graph = nx.karate_club_graph()
21# Graph = nx.from_numpy_array(np.genfromtxt("A.csv", delimiter=’,’))
22
23## PARAMETERS
24num_runs = 20
25resolution = 1 #
26
27## Initiate the quantum processor and generate the QUBO
28adv_sampler = AdvantageSampler(
29 Graph,
30 resolution=resolution,
31 num_reads=100,
32 use_clique_embedding=True
33)
34adv_iterative= IterativeSearcher(adv_sampler)
35communities, modularities, times = adv_iterative.run(num_runs=num_runs, save_results=False)
36
37# Only the community with the highest modularity
38community_structure = communities[modularities.argmax()]
39modularity = modularities.max()
40elapsed_time = times[modularities.argmax()]
Listing 1: Modularity maximization and community detection using Qommunity. After importing the relevant Python modules maximizing the modularity of the network using quantum computing would be trivial.
1import networkx as nx
2import numpy as np
3import matplotlib.pylab as plt
4
5## IMPORT THE RELEVANT MODULES
6from Qommunity.samplers.hierarchical.advantage_sampler import AdvantageSampler
7from iterative_searcher.iterative_searcher import IterativeSearcher
8from dendro import Dendrogram
9
10## PARAMETERS
11num_runs = 20
12resolution = 1
13
14## NETWORK
15G = nx.karate_club_graph()
16# Graph = nx.from_numpy_array(np.genfromtxt("A.csv", delimiter=’,’))
17
18# COMMUNITY DETECTION USING RECUSRIVE QUANTUM ANNEALING
19searcher = IterativeSearcher(
20 AdvantageSampler(
21 G,
22 resolution=resolution,
23 num_reads=100,
24 use_clique_embedding=True
25 )
26)
27results = searcher.run_with_sampleset_info(num_runs=num_runs, save_results=False, saving_path=None, iterative_verbosity=0)
28
29## MODULARITY MAXIMIZATION (Algorithm 2 in the main Methods)
30mods_Adv = results.modularity
31communities = results.communities[mods_Adv.argmax()]
32division_tree = results.division_tree[mods_Adv.argmax()]
33division_modularities = results.division_modularities[mods_Adv.argmax()]
34time = results.time[mods_Adv.argmax()]
35
36## PLOTTING AND CUSTOMIZING THE HIERARCHY PLOT
37dendro = Dendrogram(G, communities, division_modularities, division_tree)
38fig, ax = plt.subplots(1,1,figsize=(13,6))
39dendro.draw(
40 display_leafs=False,
41 yaxis_abs_log=True,
42 ax=ax,
43 fig=fig,
44 communities_labels=["Group 1", "Group 2", "Group 3", "Group 4", "Group 5", "Group 6", "Group 7", "Group 8"],
45 fig_saving_path="./Karate/dendrogram.svg",
46 title=’Hierarchicies inside the Karate network’
47)
Listing 2: Example to obtain the hierarchical structure of a given network using modularity maximization.